OpenStudy (maddiie):

Find an equation of the line satisfying the conditions given. Express your answer in standard form. Parallel to 3x-y=-5 and passing through(1-,0)

OpenStudy (anonymous):

3x-y=3

OpenStudy (radar):

Do you know what the slope is for your line?

OpenStudy (anonymous):

if the given is (-1,0) then 3x-y=-3

OpenStudy (radar):

Express you equation in the slope intercept form y=3x+5. The slope is the coefficient of x or a +3. The slope of the parallel line must be +3 also. This is the rule.

OpenStudy (radar):

You have been given one point for this parallel, that is the point (-1,0). Or I think you have....that point looks funny with the - stuck between the x and y position like that. From your study of the line equation,you know slope or m=\[\Delta y \over \Delta x\]

OpenStudy (radar):

so 3= (y-0)/(x+1) using the algebra you have learned this then becomes y=3x+3 but your problem stated it wanted it expressed in standard form which would be: 3x-y+3=0 for the standard Ax+By+C=0

OpenStudy (radar):

Hope this helps, good luck in your pursuit of knowledge.